# FedMKT: Federated Mutual Knowledge Transfer for Large and Small Language Models

### v1

@@ 4.2 Heterogeneous Setting: Performance Evaluations
As illustrated in Table 3, in comparison to both LLM-ZS and Centralized-LLM, FedMKT-LLM surpasses the zero-shot performance of LLMs and rivals the fine-tuning performance of Centralized-LLM. For instance, in the RTE QA dataset, FedMKT-LLM outperforms LLM-ZS by 30% and achieves a performance level that is nearly on par with Centralized-LLM, reaching approximately 96% of its fine-tuning performance. Similarly, in the DialogSum instruction-following dataset, FedMKT-LLM outperforms LLM-ZS by a significant 185% and achieves approximately 90% of the fine-tuning performance of Centralized-LLM. This significant achievement signifies that FedMKT-LLM effectively facilitates the acquisition of knowledge from all clients by the server. This process enables the LLM to be fine-tuned utilizing client data without necessitating access to the entire model. Moreover, the clients’ raw private data is not required to be uploaded to the LLM server, thus preserving privacy while enabling the LLM to leverage the collective knowledge of all clients.
Task Method LLaMa2-7B
RTE LLM-ZS 63.1
Centralized-LLM 85.9
FedMKT-LLM 82.3
WIC LLM-ZS 50.3
Centralized-LLM 70.4
FedMKT-LLM 61.3
BoolQ LLM-ZS 70.1
Centralized-LLM 87.6
FedMKT-LLM 84.6
CQA LLM-ZS 39.5
Centralized-LLM 69.5
FedMKT-LLM 71.8
OpenBookQA LLM-FS 31.8
Centralized-LLM 39.2
FedMKT-LLM 37.8
ARC-E LLM-ZS 69.3
Centralized-LLM 76.9
FedMKT-LLM 76.1
ARC-C LLM-ZS 40.0
Centralized-LLM 48.9
FedMKT-LLM 44.7
S-NI LLM-ZS 12.0
Centralized-LLM 49.3
FedMKT-LLM 41.4
DialogSum LLM-ZS 8.5
Centralized-LLM 27.7
FedMKT-LLM 24.2
Table 3: Heterogeneous Setting LLM Performance Comparison: FedMKT versus Baseline.
4.3 Homogeneous Setting: Performance Evaluations
In the homogeneous setting, we conduct experiments with two different model configurations, as shown in Table 1. These configurations include a combination of server-side LLaMa2-7B with client-side LLaMa2-1.3B, and another configuration pairing server-side LLaMa2-7B with client-side OPT-1.3B. It’s important to note that the number of clients is four, and within each client, all the SLMs is homogeneous. In this setting, to facilitate a comparison with FedAvg, we incorporated a FedAvg process involving all four clients within the FedMTK-SLM framework subsequent to mutual knowledge transfer. Furthermore, it’s worth noting that the standalone score here represents the average score obtained from the four clients working independently.
Tables 4 demonstrate the performance comparisons of our approach against other baselines in the homogeneous setting. The performance comparisons encompass both LLM and SLMs. In this setting, FedMTK-SLM exhibits better performance than the Standalone and FedAvg approaches due to the assistance provided by LLM. Meanwhile, in comparison to both LLM-ZS and Centralized-LLM, FedMKT-LLM surpasses the zero-shot performance of LLMs and rivals the fine-tuning performance of Centralized-LLM.
Task Method LLaMa2 OPT
CQA SLM-ZS 30.1 41.9
@@ 4.3 Homogeneous Setting: Performance Evaluations
FedMKT-SLM-FedAvg 58.6 61.5
LLM-ZS 39.5 39.5
Centralized-LLM 69.5 69.5
FedMKT-LLM-FedAvg 68.8 71.3
ARC-C SLM-ZS 26.7 23.4
@@
FedMKT-SLM-FedAvg 31.7 29.9
LLM-ZS 40.0 40.0
Centralized-LLM 49.4 49.4
FedMKT-LLM-FedAvg 46.2 46.2
ARC-E SLM-ZS 53.1 57.0
Standalone 60.3 58.8
FedAvg 60.6 59.0
FedMKT-SLM-FedAvg 61.7 60.0
LLM-ZS 69.3 69.3
Centralized-LLM 75.5 76.9
FedMKT-LLM-FedAvg 74.9 75.8
Table 4: Homogeneous Setting Performance Comparison: FedMKT-trained LLM & SLMs versus Baseline-trained LLM & SLMs.
4.4 One-to-One Setting: Performance Evaluations
In the One-to-One setting, we conduct experiments with two different model configurations, as shown in Table 1. These configurations include a combination of server-side LLaMa2-7B with client-side LLaMa2-1.3B, and another configuration pairing server-side LLaMa2-7B with client-side OPT-1.3B. Noth that both configurations involve only one client.
Tables 5 demonstrate the performance comparisons of our approach against other baselines in the one-to-one setting. The performance comparisons encompass both LLM and SLMs. In this setting, FedMTK-SLM exhibits better performance compared to both the Standalone and LLM2SLM approaches. Compared to Standalone, FedMTK-SLM achieves superior results due to the assistance provided by LLM. When compared to LLM2SLM, FedMTK-SLM further improves performance by updating the LLM, which incorporates knowledge from client-side SLM. These findings underscore the benefits of federated mutual knowledge transfer between LLMs and SLMs. Meanwhile, compared to both LLM-ZS and Centralized-LLM, FedMKT-LLM surpasses LLMs’ zero-shot performance and rivals Centralized-LLM’s fine-tuning performance.
Task Method LLaMa2 OPT
RTE SLM-ZS 49.8 52.7
Standalone 55.6 62.5
LLM2SLM 57.0 64.3
FedMKT-SLM 61.0 63.2
LLM-ZS 63.1 63.1
Centralized-LLM 85.9 85.9
FedMKT-LLM 82.3 79.1
CQA SLM-ZS 30.1 41.9
@@ 4.4 One-to-One Setting: Performance Evaluations
FedMKT-SLM 56.84 60.7
LLM-ZS 39.5 39.5
Centralized-LLM 69.5 69.5
FedMKT-LLM 69.0 71.0
ARC-C SLM-ZS 26.7 23.4
@@
FedMKT-SLM 30.8 30.4
LLM-ZS 40.0 40.0
Centralized-LLM 49.4 49.4
FedMKT-LLM 45.9 45.8
ARC-E SLM-ZS 53.1 57.0
@@
LLM2SLM 60.7 59.1
FedMKT-SLM 60.8 59.7
LLM-ZS 69.3 69.3
Centralized-LLM 75.5 76.1
FedMKT-LLM 74.8 75.8
Table 5: One-to-One Setting Performance Comparison: FedMKT-trained LLM & SLMs versus Baseline-trained LLM & SLMs.
5 Conclusions and Future Work
@@
In this work, we have presented FedMKT, a parameter-efficient federated mutual knowledge transfer framework tailored for large and small language models. FedMKT bridges the gap between the server-side large language model (LLM) and clients’ small language models (SLMs), enabling bidirectional knowledge transfer while preserving data privacy. Through extensive experiments across three distinct scenarios — Heterogeneous, Homogeneous, and One-to-One — we have demonstrated the effectiveness of FedMKT By utilizing a range of publicly available LLMs and SLMs on various NLP text generation tasks, we’ve shown that FedMKT simultaneously boosts the performance of both LLMs and SLMs. The results highlight the framework’s adaptability and practicality, paving the way for integrating LLMs into real-world applications while protecting data privacy. As large language models continue to transform the landscape of artificial intelligence, FedMKT offers a promising approach for their responsible and effective utilization.
Limitations and Future works. Our current study is constrained by computational and storage resources, preventing us from exploring larger model sizes. We plan to address this in future research.

### v2

@@ 4.2 Evaluation on Heterogeneous Setting
Table 2 also shows that FedMKT outperforms Zero-Shot and Centralized on the LLaMa2-7B of the server. For instance, on the RTE QA dataset, FedMKT outperforms Zero-Shot by 30% and achieves a performance level that is nearly on par with Centralized, reaching approximately 96% of its fine-tuning performance. This significant achievement signifies that FedMKT effectively facilitates the acquisition of knowledge from all clients by the server.
4.3 Evaluation on Homogeneous Setting
We conduct experiments with two Homogeneous settings, as shown in Table 1. The first setting (denoted as S1) involves one server-side LLaMa2-7B and four client-side LLaMa2-1.3B. The second setting (denoted as S2) involves one server-side LLaMa2-7B and four client-side OPT-1.3B.
Table 3 reports the performance comparisons of FedMKT against baselines in the two Homogeneous settings. The top sub-table and the bottom sub-table compare the performance of FedMKT against baselines on the server’s LLM and clients’ SLMs, respectively.
The top sub-table of Table 3 shows that FedMKT significantly outperforms Zero-Shot on the server’s LLM (i.e., LLaMa2-7B) in the two Homogeneous settings. It also shows that FedMKT achieves comparable performance of the Centralized scenario, in which the server’ LLM is fine-tuned using all clients’ data and the public data combined.
The bottom sub-table of Table 3 shows that FedMTK performs better than the Zero-Shot, Standalone, and FedAvg due to the assistance of the server’s LLM. For example, in the CQA dataset, FedMKT outperforms FedAvg by 4% on the LLaMa2-1.3 SLM and by 5% on the OPT-1.3B SLM, respectively.
Task Method S1: Server LLaMa2-7B S2: Server LLaMa2-7B
CQA Zero-Shot 39.5 39.5
Centralized 69.5 69.5
FedMKT 68.8 71.3
ARC-C Zero-Shot 40.0 40.0
Centralized 49.4 49.4
FedMKT 46.2 46.2
ARC-E Zero-Shot 69.3 69.3
Centralized 75.5 75.5
FedMKT 74.9 74.8
Task Method S1: Clients LLaMa2-1.3B S2: Clients OPT-1.3B
CQA Zero-Shot 30.1 41.9
@@ 4.3 Evaluation on Homogeneous Setting
FedMKT 58.6 61.5
ARC-C Zero-Shot 26.7 23.4
@@
FedMKT 31.7 29.9
ARC-E Zero-Shot 53.1 57.0
Standalone 60.3 57.9
FedAvg 60.6 58.8
FedMKT 61.7 60.1
Table 3: Method Performance Comparison in Homogeneous settings. We evaluate FedMKT using two homogeneous settings. The first setting (denoted as S1) involves one server-side LLaMa2-7B LLM and four client-side LLaMa2-1.3B SLMs, while the second setting (denoted as S2) involves one server-side LLaMa2-7B LLM and four client-side OPT-1.3B SLMs. The top and bottom sub-tables compare the performance of FedMKT against baselines on the server’s LLM and clients’ SLMs, respectively. The results reported in the bottom sub-table are the average of all clients.
4.4 Evaluation on One-to-One Setting
We evaluate FedMKT using two One-to-One settings. The first setting (denoted as S1) involves one server-side LLaMa2-7B LLM and one client-side LLaMa2-1.3B SLM, while the second setting (denoted as S2) involves one server-side LLaMa2-7B LLM and one client-side OPT-1.3B SLM.
Table 4 reports the performance comparisons of FedMKT against baselines in the two One-to-One settings. The top and bottom sub-tables compare the performance of FedMKT against baselines on the server’s LLM and clients’ SLMs, respectively.
The top sub-table of Table 4 shows that FedMKT notably surpasses Zero-Shot and rivals Centralized on the performance of the server’s LLM. The bottom sub-table of Table 4 shows that FedMTK achieves superior SLM performance over Zero-Shot, Standalone, and LLM2SLM due to the assistance of LLM. These empirical results demonstrate the effectiveness of FedMKT in transferring knowledge between the LLM and SLMs.
Task Method S1: Server LLaMa2-7B S2: Server LLaMa2-7B
CQA Zero-Shot 39.5 39.5
Centralized 69.0 68.3
FedMKT 69.0 71.0
ARC-C Zero-Shot 40.0 40.0
Centralized 45.9 48.6
FedMKT 45.9 45.8
ARC-E Zero-Shot 69.3 69.3
Centralized 74.4 73.6
FedMKT 74.8 74.8
Task Method S1: Clients LLaMa2-1.3B S2: Clients OPT-1.3B
CQA Zero-Shot 30.1 41.9
@@ 4.4 Evaluation on One-to-One Setting
FedMKT 56.84 60.7
ARC-C Zero-Shot 26.7 23.4
@@
FedMKT 30.8 30.4
ARC-E Zero-Shot 53.1 57.0
@@
LLM2SLM 60.7 58.4
FedMKT 60.8 58.5
Table 4: Method Performance Comparison in One-to-One settings. We evaluate FedMKT using two one-to-one settings. The first setting (denoted as S1) involves one server-side LLaMa2-7B LLM and one client-side LLaMa2-1.3B SLM, while the second setting (denoted as S2) involves one server-side LLaMa2-7B LLM and one client-side OPT-1.3B SLM. The top and bottom sub-tables compare the performance of FedMKT against baselines on the server’s LLM and a client’s SLM, respectively.
5 Conclusions
@@
In this study, we have presented FedMKT, a parameter-efficient federated mutual knowledge transfer framework tailored for large and small language models. FedMKT bridges the gap between the server-side LLM and clients’ SLM, enabling selective mutual knowledge transfer while preserving data privacy. Through extensive experiments across three distinct scenarios, we have demonstrated that FedMKT simultaneously boosts the performance of both LLMs and SLMs.
Limitations
In this study, we transfer knowledge between the server and clients using logits of a public dataset, motivated by efficiency and privacy considerations. Although empirical evidence suggests that sharing logits of public datasets between the server and clients is more privacy-preserving than sharing model gradients or parameters Li and Wang (2019); Cho et al. (2022), there is no theoretical guarantee that this approach does not compromise the privacy of clients’ sensitive data. This issue warrants further investigation. Furthermore, our study is limited by computational and storage constraints, which preclude the exploration of larger language models. This highlights the inherent trade-off between utility and efficiency. Our future research aims to investigate and optimize this trade-off.
Appendix A FedMKT Workflow
The workflow of FedMKT is outlined as follows:
  1. 1.
  2. 2.
    After local training, each client $k$ computes a knowledge set $\mathcal{S}_{k}=\{l^{i}_{k},p^{i}_{k}\}_{i=1}^{N}$ consisting of loss-logit pairs through its local model based on the public dataset.
  3. 3.
    Each client $k$ upload $\mathcal{S}_{k}$ to the server.
  4. 4.
  5. 5.
    On the server side, knowledge is selected from the SLMs to the LLM according to Algorithm 2.
  6. 6.
    On the server side, knowledge is transferred from the SLMs to the LLM based on the selected knowledge.
  7. 7.
    Once the knowledge transfer from SLMs to LLM is completed on the server, the server then computes a knowledge set $\mathcal{S}_{0}=\{l^{i}_{0},p^{i}_{0}\}_{i=1}^{N}$ consisting of loss-logit pairs through LLM based on the public dataset.
  8. 8.
    The server disseminates $\mathcal{S}_{0}$ to all the clients.
  9. 9.
  10. 10.
    On the client side, knowledge is selected from the LLM to each client SLM according to Algorithm 2.
  11. 11.
    On the client side, knowledge is transferred from the LLM to each client SLM based on the selected knowledge.
Appendix B Implementation Details of Token Alignment
In our work, we engage in a bidirectional token alignment procedure, encompassing the alignment of SLM tokens with their corresponding LLM tokens, and vice versa. Both alignments adhere to a similar methodology. Presently, we shall elaborate on the process of aligning LLM tokens with their matching SLM tokens. To map the predicted token logits from the LLaMa2-7B (LLM) model to the Bloom-1.1B (SLM) model, several steps must be undertaken. The detailed process is as follows:
  1. 1.
    Building an Optimal Vocabulary Mapping Table:
    1. (a)
      For each token in the LLaMa2 vocabulary, iterate through the Bloom vocabulary.
    2. (b)
      Use edit distance as a similarity measure to find the closest token in the Bloom vocabulary to the token in the LLaMa2 vocabulary.
    3. (c)
      If there are multiple token with the same minimum edit distance, choose the one with the lexicographically smallest order.
    4. (d)
      Save this mapping relationship in the optimal vocabulary mapping table.
  2. 2.
    Tokenization and Alignment:
    1. (a)
      Tokenize the sentence "we utilize the dynamic programming approach to align tokens" using both the LLaMa2 and Bloom tokenizers.
    2. (b)
      To align the two tokenization results and determine the optimal matching path, we utilize a dynamic programming algorithm. As an illustration, consider the tokenization outputs from LLaMa2 and Bloom. LLaMa2’s tokenization yields: [’we’, ’util’, ’ize’, ’the’, ’dynamic’, ’programming’, ’approach’, ’to’, ’align’, ’tokens’]. In contrast, Bloom’s tokenization produces: [’we’, ’utilize’, ’the’, ’dynamic’, ’programming’, ’approach’, ’to’, ’align’, ’tokens’]. In this instance, seven terms from LLaMa2 align perfectly with those from Bloom, such as "we" and "dynamic". Notably, the LLaMa2 tokens ’util’ and ’ize’ collectively map to the single Bloom token ’utilize’. In scenarios where multiple tokens align to one, like the 2-to-1 case of ’util’ and ’ize’ mapping to ’utilize’, we consider ’utilize’ as a match for ’util’ based on an optimal vocabulary mapping.
  3. 3.
    Logits Mapping:
    1. (a)
      Iterate through each token $t_{t}$ in the Bloom tokenization result.
    2. (b)
      For each $t_{t}$, check if it uniquely matches a token $t_{s}$ in the LLaMa2 tokenization result.
    3. (c)
      If $t_{t}$ uniquely matches $t_{s}$, then for each token $t_{p}$ in the Top-$K$ predicted token of $t_{s}$ from LLaMa2 and its corresponding logit $logit_{p}$: Find the position $pos$ in the Bloom vocabulary that corresponds to $t_{p}$ using the optimal vocabulary mapping table. If $pos$ has not been assigned a value before, copy $logit_{p}$ to the corresponding position in the Bloom logits distribution matrix $logit_{t}$.
    4. (d)
      If $t_{t}$ does not have a unique match, generate one-hot logits for $t_{t}$.
  4. 4.
    Processing the Results:
    1. (a)
      Ultimately, each token $t_{t}$ in Bloom will have a corresponding logits distribution matrix $logit_{t}$.
    2. (b)
      These logits can be directly used for subsequent training in the Bloom model.
Appendix C Computation and Communication Complexity
Appendix D More on Experimental Details
D.1 Hyperparameter Settings
LoRA Parameters. We utilized the PEFTMangrulkar et al. (2022) library with the following configurations: r=8, lora_alpha=16, lora_dropout=0.05.
Common Parameters for LLM and SLMs. We set batch_size=4, used the AdamW optimizer with adam_beta1=0.9 and adam_beta2=0.95. The warmup_ratio was set to 0.008, the weight_decay was 0.1, max_grad_norm was 1.0. The $\lambda$ was 0.9. The number of training rounds for all data is within 10 and the number of training rounds for different datasets may be different.
LLM Parameters. During distillation, the local epoch R was set to 1. The learning rates $\eta_{\omega}$ were specified as 3e-5 for the datasets RTE/WIC/BoolQ/CQA/ARC-C/DialogSum/S-NI, and 2e-5 for ARC-E.
SLM Parameters. During training for the four clients, the local epoch E was set to 1. The learning rates $\eta_{\theta}$ were as follows: for "OPT-1.3b", $\eta_{\theta}$=3e-5; for "GPT-2-xlarge", $\eta_{\theta}$=3e-4; for "Bloom-1b1", $\eta_{\theta}$=3e-5; and for "LLaMa-2-1.3b", the same learning rates as for the LLM were used.
D.2 Data Splitting
For the datasets RTE/WIC/BoolQ/CQA/ARC-E/ARC-C/DialogSum, we randomly split the training data into five equal parts, with one part serving as the public dataset and the remaining four parts as private dataset for the four clients. All these datasets(including train, validate, test) were downloaded from HuggingFaceLhoest et al. (2021). For the S-NI dataset, we first processed the data using minillmGu et al. (2023) to retain samples with an output length greater than or equal to 11. From this processed data, we randomly selected 300 samples as the evaluation dataset. The remaining data was then split into five equal parts, with one part serving as the public dataset and the other four parts as private data for the four clients.
D.3 Dataset Licenses
For the datasets RTE/WIC/BoolQ/CQA/ARC-E/ARC-C/DialogSum were downloaded from HuggingFaceLhoest et al. (2021) and under Apache License, Version 2.0. For the S-NI dataset, it was from minillmGu et al. (2023) and under MIT License.
D.4 Machine Configuration
The experiments were conducted on machines equipped with either 4 Nvidia V100 32G or 8 Nvidia V100 32G GPUs.